OpenStudy (anonymous):

The surface areas of two similar solids are 321 yd2 and 1117 yd2. The volume of the larger solid is 1772 yd3. What is the volume of the smaller solid? A) 1,772 yd3 B) 950 yd3 C) 764 yd3 D) 273 yd3 Thanks!

OpenStudy (saifoo.khan):

its easy though

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

k one sec

OpenStudy (anonymous):

17.9

OpenStudy (anonymous):

33.4

OpenStudy (anonymous):

what now?

OpenStudy (saifoo.khan):

im getting 268, wait. let me C

OpenStudy (anonymous):

Any luck?

OpenStudy (saifoo.khan):

yay!

OpenStudy (saifoo.khan):

272.98

OpenStudy (saifoo.khan):

D

OpenStudy (anonymous):

Thanks!

OpenStudy (saifoo.khan):

[:

OpenStudy (anonymous):

9mathkid4747: "The surface areas of two similar solids are 321 yd2 and 1117 yd2. The volume of the larger solid is 1772 yd3. What is the volume of the smal" Let d, the diameter of a solid, be the straight-line distance between the two farthermost points on its surface. If S is its surface area and V is its volume, then there positive real numbers h and k such that S = h*d^2, and V = k*d^3. Let the smaller body have diameter d1, surface area S1, and volume V1; and let the larger body have diameter d2, surface area S2, and volume V2. Then S1 = h*d1^2, S2 = h*d2^2, V1 = k*d1^3, and V2 = k^d2^2. The area equations reveal that S1/S3 = d1^2/d2^2, and the volume equations reveal that V1/V2 = d1^3/d2^3. Solving these last two equations fot d1/d2 gives us d1/d2 = (S1/S2)^(1/2) and d1/d2 = (V1/V3)^(1/3), so now we can set (S1/S2)^(1/2) = (V1/V2)^(1/3) and solve for V1 = V2*(S1/S2)^(3/2). = 1772*(321/1117)*(3/2) = 273 yd^3 approximately.